OpenStudy (ramennoodle01):

Solve this equation using elimination by addition: 4m-n=13 m-4n=-23

OpenStudy (triciaal):

the objective is to make the coefficient of your chosen variable equal and opposite so when you add both equivalent equations that variable will be eliminated. you now have an equation with one unknown. solve to get the value of this variable. now substitute this value in either of the original and get the value of the variable that was eliminated. it is always a good practice to check both values in either of the originals because it is very easy to make arithmetic errors.

OpenStudy (ramennoodle01):

I figured it out. Thank you! The answer I got was 5 and 7

OpenStudy (triciaal):

you need m = and n =

OpenStudy (ramennoodle01):

M=5 and N=7 I did -4(4m-n)=13(-4) 1(m-4n)=-23(1)

OpenStudy (triciaal):